Jeremy baked a cake. Sue cut the cake into equal-sized pieces and took one piece. Nathan took 1/5 of the remaining pieces. Finally, Jeremy took 1/3 of the remaining pieces. If 1⁄2 of the cake is still remaining, how many pieces did Sue cut the cake into?
Answer by Boreal(15235)

piece size=1/x
cut into x pieces
1-(1/x) after Sue
(1-(1/x))-(1-(1/x)/5) after Nathan.
This is (x-1)/x - [(x-1)/5] after Nathan.
this is (4/5)(x-1)/x
Now, 4(x-1)/5x is left
After Jeremy takes 1/3, there is 2/3 of that left or 8(x-1)/15x left.
That equals 1/2
16(x-1)/15x=1
16x-16=15x
x=16 pieces ANSWER
after Sue, 15/16 is left
Nathan took 1/5 of that so that 12/16 pieces are left
Jeremy took 1/3 of that so now 8/16 pieces are left or half.